OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 2x + y = −4 5x + 3y = −6

OpenStudy (anonymous):

It's -4 right?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (sleepyjess):

have you learned about solving with substitution?

OpenStudy (sleepyjess):

@ilikemath50

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

ok that makes things easier to explain :-) do you know how to start this?

OpenStudy (sleepyjess):

if not that is fine

OpenStudy (anonymous):

Yes I have, but I am sort of forgetting how to..

OpenStudy (sleepyjess):

ok since we are solving for y we want the y on the first equation by itself. Can you do that?

OpenStudy (anonymous):

2x + y = - 4 y = -2x - 4

OpenStudy (anonymous):

then we substitute -2x - 4 in for y for the equation 5x + 3y = −6 right?

OpenStudy (anonymous):

5x + 3(-2x - 4) = -6

OpenStudy (anonymous):

am i doing okay so far? I don't remember this very well so im doing my best lol

OpenStudy (anonymous):

5x - 6x - 12 = - 6

OpenStudy (anonymous):

-x = -6 + 12 -x = 6 or x = -6 and we substitute x for -6 in the equation 2x + y = - 4

OpenStudy (anonymous):

2(-6) + y = -4 -12 + y = -4

OpenStudy (anonymous):

y = -4 + 12 y = 8 then we check the answer...

OpenStudy (anonymous):

2x + y = -4 2(-6) + 8 = -4 -12 + 8 = -4 -4 = -4 y = -4

OpenStudy (anonymous):

done

OpenStudy (anonymous):

I'm doubting myself a lil bit though, lol

OpenStudy (anonymous):

are you there? @sleepyjess